# Transaction Profiling and Address Role Inference in Tokenized U.S. Treasuries

### v1

@@ 5.2.4 Experiment Result
The results in Table 7 compare the performance of benchmark methods and our proposed PoincaVec variants across four blockchain transaction graph datasets with varying structural properties and label sparsity. Across all datasets, the full PoincaVec pipeline incorporating both hierarchical radius-based features (H) and topology-aware walk embeddings (T) achieves the highest or near-highest F1 scores, demonstrating strong generalization to fraud node (address) classification tasks.
@@
For the Ethereum dataset featuring a large, sparsely labeled transaction network, PoincaVec (w/ H, w/ T) achieves an F1 of 0.940, outperforming proximity-based baselines such as Node2Vec and structurally driven Role2Vec. The AscendEXHacker dataset exhibits an imbalanced role distribution, where only a small fraction of addresses are tagged as direct heist participants, while the majority are exchange-facing services such as DEX routers (e.g., Uniswap). In this setting, hierarchical radius-based features (H) offer an inductive prior: heist addresses tend to lie peripherally with sparse connectivity, while services cluster near radius enabling PoincaVec (w/ H, w/o T) to achieve the highest F1 of 0.840, whereas adding walk-based features reduces performance to 0.794. Notably, PoincaVec obtained an F1 of 0.945 on the Ethereum Classic (ETC) dataset, demonstrating its capacity to disentangle abnormal node roles. These results demonstrate the performance of embedding blockchain transaction graphs in hyperbolic space, where latent hierarchies and peripheral node behaviours, which are common in fraud and scam context, can be captured more compactly in our proposed method.
6 Conclusion
@@
This study presents the first transaction-level analysis of tokenized U.S. Treasuries across multiple blockchain networks. By decoding contract interactions and profiling function usage, we identify dominant operational roles and reveal distinct patterns between institutional and retail participants. Our cross-chain comparison highlights how token design and network context shape transactional behaviour. Furthermore, we introduce a hyperbolic embedding model that infers address roles. These findings provide a quantitative account of transaction behaviours across tokenized U.S. Treasuries, detailing how operational functions and participant roles emerge across chains and transactional contexts. The proposed Poincaré-embedding-based model enables role identification and achieves improved performance both on our RWA dataset and in generalizing to external classification and anomaly detection tasks.
7 Appendix: Hyperbolic Radius as a Proxy for Latent Hierarchy
$r_{h}:=\|\phi(u)\|=\tanh\left(\frac{h\cdot\ell}{2}\right),$
where $\ell>0$ is the fixed geodesic step length in the hyperbolic metric. Then:
Lemma 7.1.
@@ 7 Appendix: Hyperbolic Radius as a Proxy for Latent Hierarchy
The hyperbolic radius $r_{h}$ is a strictly increasing function of tree depth $h$.
Proof 7.2.
@@
Therefore, $r_{h}$ increases monotonically with $h$.
Corollary 7.3.
Under tree-consistent Poincaré embeddings, node radius $\|\mathbf{z}_{v}\|$ can be interpreted as a continuous proxy for latent hierarchical depth.

### v2

@@ V-B6 Extended Evaluation on Public Blockchain Transaction Datasets
Table VII presents a comparison of benchmark methods and our PoincaVec variants across four blockchain transaction graph datasets with varying structural properties and labels. Across all datasets, the full PoincaVec pipeline incorporating both hierarchical radius-based features (H) and topology-aware walk embeddings (T) achieves the highest or near-highest F1 scores, demonstrating generalization to anomalous address classification tasks.
@@
For the large, sparsely labeled network of Ethereum dataset, PoincaVec (w/ H, w/ T) achieves an F1 of 0.940, outperforming proximity-based baselines such as Node2Vec and structurally driven Role2Vec. The AscendEXHacker dataset exhibits an imbalanced role distribution, where only a small fraction of addresses are tagged as direct heist participants, while the majority are exchange-facing services such as DEX routers (e.g., Uniswap). In this setting, hierarchical radius-based features (H) offer an inductive prior: heist addresses tend to lie peripherally with sparse connectivity, while services cluster near radius, enabling PoincaVec (w/ H, w/o T) to achieve the highest F1 of 0.840, whereas adding walk-based features reduces performance to 0.794. Notably, PoincaVec obtained an F1 of 0.945 on the Ethereum Classic (ETC) dataset, demonstrating its capacity to disentangle abnormal node roles. These results demonstrate the performance of embedding blockchain transaction graphs in hyperbolic space, where latent hierarchies and peripheral node behaviours, which are common in fraud and scam contexts, can be captured more compactly in our proposed method. Additional evaluation figures and detailed metric analyses are provided in Appendix D, including full Precision@k curves and prevalence-sensitivity visualizations across all datasets.
VI Conclusion
@@
This study presents the first transaction-level analysis of tokenized U.S. Treasuries across multiple blockchain networks. By decoding contract function usage, we identify dominant operational roles and reveal distinct patterns between institutional and retail participants, and provide a quantitative account of how institutional and retail participants engage with tokenized U.S. Treasuries across chains and transactional contexts. We proposed a Poincaré embedding-based model that enables role inference and demonstrates improved performance both on our RWA dataset and when generalizing to external anomalous address classification tasks.
Appendix A Datasets details
Ethereum Transaction [52] An Ethereum transaction dataset constructed from historical blockchain records, containing 2,973,489 addresses and 13,551,303 edges. A total of 1,165 addresses are labeled as illicit based on curated forensic data, while the rest remain unlabeled. This dataset has an average node degree of 4.56.
AscendEXHacker [51] A network extracted via Etherscan’s Heist label, focusing on the 2021 AscendEX exchange exploit. It contains 6,642 addresses across 29,074 transactions. Ground-truth annotations label 84 addresses as direct heist participants, while 638 addresses are identified as DEXs or Uniswap-related service accounts.
PlusTokenPonzi [51] A transaction dataset linked to the PlusToken Ponzi, one of the largest crypto scams. It comprises 34,521 unique accounts and 58,049 transactions, of which 30,782 addresses are explicitly identified as scam participants.
Ethereum Classic [3] A labeled transaction network from the Ethereum Classic (ETC) network, compiled using EtherscamDB scam reports. The dataset comprises 73,034 nodes and 71,250 edges, with 169 addresses labeled as scammers based on crowdsourced and externally verified annotations.
Appendix B Hierarchical Radius-based Features Vector
$r_{v}=2\cdot\tanh^{-1}(\|\mathbf{z}_{v}\|),$
which reflects the geodesic distance from the origin and serves as a proxy for depth in the hierarchy, i.e., nodes closer to the origin are higher in the hierarchy, while those near the boundary are deeper.
From the neighbour radii, we compute an 11-dimensional hierarchical feature vector $\mathbf{h}_{v}\in\mathbb{R}^{11}$ comprising the following components:
$r_{v}$ Absolute depth (node’s self radius)
$\mathbf{b}_{v}$ 4-bin histogram of relative radii $(r_{u}-r_{v})$
over range [1,1][-1,1]
Altogether, the hierarchical feature vector is given by:
Appendix C Hyperbolic Depth Derivation
@@
Lemma. The hyperbolic radius $r_{h}$ is a strictly increasing function of tree depth $h$.
Proof.
@@
Therefore, $r_{h}$ increases monotonically with $h$. ∎
Corollary. Under tree-consistent Poincaré embeddings, node radius $\|\mathbf{z}_{v}\|$ can be interpreted as a continuous proxy for latent hierarchical depth.
Appendix D Extended Evaluation Metrics
D-1 Evaluation of Ranking Performance via Precision@k
To assess ranking fidelity, we evaluate Precision@k over a range of thresholds $k\in\{\,0.05\cdot n\mid n=1,...,12\,\}\cdot|\text{test set}|$, defined as fractions of the test set size. We evaluate top-$k$ ranking performance using Precision@k on all five datasets. As shown in Figure 5, our proposed PoincaVec variants consistently rank positive instances more effectively than baselines across a range of thresholds. In datasets such as RWA and AscendEXHacker, PoincaVec (w/ H, w/ T) delivers substantial gains in early precision, confirming its suitability for skewed detection settings. On denser class datasets such as PlusTokenPonzi, where nearly all methods saturate near-perfect precision, the performance of the proposed method remains stable and competitive. These results suggest that PoincaVec inductively improves F1-oriented retrieval without compromising ranking precision.
D-2 Prevalence-Sensitivity of Performance Improvements
To evaluate the effect of positive class prevalence, we assess prevalence-sensitivity by relating the performance improvement of each dataset to the conducted experiment. Figure 6 summarizes two aspects: the paired macro-$F_{1}$ gains of our method over the best baseline (left), and how those gains vary with positive prevalence (right). The left panel visualizes absolute improvements for each dataset, confirming that our proposed model consistently outperforms the baselines on both our collected RWA U.S. Treasury transaction dataset and all other public blockchain transaction datasets.
The right panel plots $\Delta F_{1}$ uplift against prevalence (log scale), revealing a concave pattern: improvements peak at intermediate prevalence and diminish at both extremes. These results indicate that our method achieves the greatest improvements when the positive class is neither too rare nor too dominant, i.e., under moderate imbalance, particularly when the baseline model struggles to separate roles in scenarios where class boundaries are less distinct and prevalence alone does not strongly determine separability.
Figure 6: Paired performance comparison and prevalence-sensitivity. Left: Macro F1 scores for the best baseline vs. PoincaVec across all five datasets. Right: Improvement in macro F1 (percent) plotted against positive prevalence. Color encodes best baseline performance to contextualize potential uplift. The dashed line shows a concave fit, with relatively more improvement at intermediate prevalence levels.
Ethical Considerations This study uses only publicly available, pseudonymous blockchain transaction data and performs no linkage to real-world identities; we neither collect nor expose personally identifiable information. All analyses are conducted on‐chain transaction graphs; results are reported in aggregate and we release no per-address classifications, limiting targeting risks. The role-inference model is probabilistic tool for on-chain addresses analyzing and should not be treated as definitive evidence of illicit behaviour. We will publish code that reproduces aggregate findings. Given these safeguards, we think that the transparency benefits for market participants and researchers in the blockchain, and web for good community.